OpenStudy (anonymous):

WHAT IS THE SOLUTION TO THE PROPORTION? 12/15 = 8/X A: 16 B. 10 C. 9 D. 12

OpenStudy (amistre64):

hmmm \[12(\frac{8}{12}) = 8\] sooo \[15(\frac{8}{12}) = x\]

OpenStudy (anonymous):

so its 12

OpenStudy (amistre64):

i dont get 12 from the math ...

OpenStudy (amistre64):

\[\frac{15(8)}{12}\] \[\frac{5.3.4.2}{4.3}=5.2\]